# Controlling Multimodal Conversational Agents with Coverage-Enhanced Latent Actions

### v1

@@ Abstract
Vision-language models are increasingly employed as multimodal conversational agents (MCAs) for diverse conversational tasks. Recently, reinforcement learning (RL) has been widely explored for adapting MCAs to various human-AI interaction scenarios. Despite showing great enhancement in generalization performance, fine-tuning MCAs via RL still faces challenges in handling the extremely large text token space. To address this, we learn a compact latent action space for RL fine-tuning instead. Specifically, we adopt the learning from observation mechanism to construct the codebook for the latent action space, where future observations are leveraged to estimate current latent actions that could further be used to reconstruct future observations. However, the scarcity of paired image-text data hinders learning a codebook with sufficient coverage. Thus, we leverage both paired image-text data and text-only data to construct the latent action space, using a cross-modal projector for transforming text embeddings into image-text embeddings. We initialize the cross-modal projector on paired image-text data, and further train it on massive text-only data with a novel cycle consistency loss to enhance its robustness. We show that our latent action based method outperforms competitive baselines on two conversation tasks across various RL algorithms.
@@ 1 Introduction
Vision-language models (VLMs) yin-2024-MLLMsurvey like Qwen-VL bai-2025-qwen3vl and GPT-4o hurst-2024-gpt4ocard are increasingly employed as multimodal conversational agents (MCAs) for various conversation tasks yao-2025-MLLMAgentsurvey. MCAs enable emotionally rich and contextually grounded dialogues based on understanding both input images and texts, and thus become particularly valuable in fields like entertainment mehta-2022-exploring, online education griol-2014-developing, and personalized assistants nguyen-2024-yo.
@@
Recently, reinforcement learning (RL) sutton-1998-reinforcement has been widely explored for adapting MCAs to diverse real-world human-AI interaction scenarios zhou-2025-reinforcedMLLM. Generally, RL algorithms frame response token generation in MCAs as a sequential decision-making process chen-2021-decision, which optimize the policy to maximize cumulative rewards through interacting with environments. Despite showing great enhancement in generalization performance chu-2025-sftRL, fine-tuning MCAs via RL still faces challenges in dealing with large exploration spaces. For instance, with token vocabulary size $|\mathcal{V}|$ and maximum response length $m$, the sampling space for RL scales exponentially as $|\mathcal{V}|^{m}$.
@@
To address the challenge of large text token space, we learn a compact latent action space for RL fine-tuning instead, following previous works jia-2025-controlling. Specifically, we adopt the learning from observation mechanism jiang-2023-efficient; ye-2025-latent to construct the codebook for the latent action space, where future observations are leveraged to estimate current latent actions that could be further used to reconstruct future observations. As a result, the action sampling space at each step is reduced from the token vocabulary size $|\mathcal{V}|$ (e.g., 152K for Qwen2.5-VL bai-2025-qwen25vl) to the latent action codebook size $|\mathcal{C}|$ (e.g., 128).
@@
Generally, the codebook has to be learned from diverse data with sufficient coverage, which is a prerequisite for effective RL exploration in latent spaces chen-2025-coverage. Note that VLMs in MCAs are typically pre-trained on paired image-text corpora $(V,T)$, which implicitly convey complementary and partially redundant information between visual and textual modalities radford-2021-clip. Unfortunately, while unpaired image collections and text corpora are abundant on the web, curating them into aligned image-text corpora remains prohibitively costly gupta-2025-better, posing a dilemma in constructing latent spaces. On one hand, using limited paired data and abundant unpaired data would introduce unimodal bias zhang-2024-UniModalBias, where a model would overly rely on one modality and ignore others. On the other hand, training the codebook solely on limited paired data may result in insufficient coverage, thereby impairing the agent’s generalization ability when handling diverse unseen conversation scenarios.
@@
In this paper, we leverage both paired image-text data $(V,T)$ and unpaired text-only data $T$ to learn the coodbook for the latent space. To improve the coverage of latent actions while avoiding potentially unimodal bias, we attempt to construct pseudo paired data $(V^{\prime},T)$ based on text-only data $T$, and use the pseudo data $(V^{\prime},T)$ and the collected data $(V,T)$ to learn the cookbook.
@@
However, training a conditional image generator $G(V|T)$ for this purpose is computationally expensive due to the high dimension nature of images pope-2021-dimension. Thus, we learn a cross-modal projector $P$ instead, which transforms an input text $e^{T}$ to an image-text pair $e^{V,T}$ in the embedding space, based on the cross-modal redundancy assumption radford-2021-clip. Concretely, for each item in the paired image-text data $(V,T)$, we compute the text embedding $e^{T}$ and image-text embedding $e^{V,T}$ using an existing encoder, and train the projector $P$ to imitate the projection between these two kinds of embeddings. To enhance the robustness of the projector $P$, we further train it on massive text-only data $T$ using a cycle consistency loss zhu-2017-unpaired. We introduce an additional projector $P^{\prime}$ that can transform image-text embedding $e^{V,T}$ back to text embedding $e^{T}$. In this way, we can optimize the projector $P$ by enforcing cycle consistency on text-only data $T$ such that $P^{\prime}(P(e^{T}))\approx e^{T}$.
@@
We evaluate our method on two conversation tasks, namely multimodal role-playing conversation dai-2025-mmrole and multimodal personalized conversation li-2025-aligning. To evaluate the generalizability of latent actions, we conduct experiments using various RL algorithms, such as GRPO shao-2024-GRPO and Dr.GRPO liu-2025-DRGRPO. We construct the latent action space using paired image-text data $(V,T)$ and text-only data $T$. The $(V,T)$ data are comprised of image-caption pairs, multimodal news articles, and multimodal Wikipedia pages, totaling 14M images and 1B text tokens. The text-only data are mainly derived from SlimPajama cerebras-2023-slimpajama, which contains 627B text tokens. Experimental results show that our method outperforms competitive baselines.
@@ 2 Preliminary
In reinforcement learning (RL), problems are framed by a Markov Decision Process (MDP) $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R}\rangle$. For VLMs, the state at step $t$ is the contextual information $s_{t}=(x^{V},x^{T_{1:t}})\in\mathcal{S}$, which includes the input image $x^{V}$ and the current token sequence $x^{T_{1:t}}$. $\mathcal{A}$ is the action space containing all possible actions $a_{t}$ at each step. $\mathcal{T}$ is the state transition function, governing the transition from $s_{t}$ to $s_{t+1}$, i.e., $P\big(s_{t+1}\mid s_{t},a_{t}\big)$. The reward function $\mathcal{R}\big(x^{T_{p+1:m}}\big)$ assigns a scalar reward to the response $x^{T_{p+1:m}}$, conditioned on the input $(x^{V},x^{T_{1:p}})$, with prompt length $p$ and maximum sequence length $m$, following common practice in RL for VLMs shen-2025-vlmR1.
@@
In traditional token-level RL, each action $a_{t}$ corresponds to selecting the next text token $x^{T_{t+1}}$ from the token vocabulary $\mathcal{V}$, i.e., $\mathcal{A}=\mathcal{V}$. While in latent action RL, at each step $t$, the policy $\pi_{\theta}(a_{t}|x^{V},x^{T_{1:t}})$ selects a latent action $a_{t}$ from a compact codebook $\mathcal{C}$, i.e., $\mathcal{A}=\mathcal{C}$. During RL exploration, the latent action policy samples a latent action at each step, ultimately yielding the terminal state $s_{m}$. During exploitation, the latent action policy is refined to maximize expected rewards using RL algorithms such as GRPO shao-2024-GRPO.
@@ 3.2 Latent Action Space Learning
Following jia-2025-controlling, we construct the latent action space using large-scale corpora in two steps. 1) inverse dynamics learning, which trains the $f_{\text{world}}$, $f_{\text{inverse}}$, and $\mathcal{C}$ in an unsupervised manner (Fig. 2 (a)); 2) policy behavior cloning, which trains the policy model $\pi_{\theta}$ to mimic the latent action $a_{t}$ inferred by $f_{\text{inverse}}$ (Fig. 2 (b)).
@@ 3.3 Latent Action Reinforcement Learning
where $R(\cdot)$ denotes the reward function.
@@ 4.1 Experimental Setup
We build the language world model, inverse dynamics model, and policy model upon the same foundation vision-language model. Specifically, we use Qwen2.5-VL-3B-Instruct and Qwen2.5-VL-7B-Instruct bai-2025-qwen25vl for main experiments. The latent action space is implemented as a codebook with size $|\mathcal{C}|=128$.
@@
During the latent action space construction stage (Section 3.2), we use a mixture of paired image-text corpora $\mathcal{D}^{VT}$ and text-only corpora $\mathcal{D}^{T}$. For $\mathcal{D}^{VT}$, we collect image-caption pairs from Conceptual-12M changpinyo-2021-Conceptual, multimodal news articles from N24News wang-2022-N24News, and multimodal Wikipedia data from WikiWeb2M burns-2023-wiki, totaling 14 million images and 1 billion text tokens. For $\mathcal{D}^{T}$, we collect text-only data mainly from the SlimPajama-627B dataset cerebras-2023-slimpajama, which contains 627 billion text tokens.
For latent action RL (Section 3.3), we evaluate our method on two downstream tasks: 1) multimodal role-playing conversation on MMRole dai-2025-mmrole, where we focus on the challenging Comment subset; we train on the in-distribution (ID) split and evaluate on ID and out-of-distribution (OOD) test sets; 2) multimodal personalized conversation on PCogAlignBench li-2025-aligning, where we train the agent on the LS1 set and evaluate on LS1 and LS2 test sets.
@@
We adopt the LLM-as-a-Judge metric to evaluate model performance, using prompt templates validated by dai-2025-mmrole; li-2025-aligning, which show high correlation with human judgments. For each sample, the LLM judge scores both the model and ground-truth responses across benchmark-specific dimensions, with scores ranging 1-10. Then, following dai-2025-mmrole, we report the ratio of the model’s average score to the ground-truth response’s average score across all evaluation dimensions. We report the mean and standard deviation across three evaluation runs.
@@
We consider two categories of baselines: 1) Non-RL baselines: the naive Prompt and supervised fine-tuning (SFT); 2) RL-based methods, where we compare two optimization strategies, token-level and latent action RL, using four algorithms: a) Group Relative Policy Optimization (GRPOshao-2024-GRPO, b) Dr. GRPO liu-2025-DRGRPO, c) Decoupled Clip and Dynamic Sampling Policy Optimization (DAPOyu-2025-dapo, and d) Beta Normalization Policy Optimization (BNPOxiao-2025-BNPO. The reward functions are kept the same for methods. Please refer to the Appendix B for more experimental details.
@@ 4.2 Main Results
To thoroughly evaluate the performance of multimodal conversational agents trained with latent actions across various fine-grained conversational dimensions, following prior work dai-2025-mmrole; li-2025-aligning, we assess eight dimensions on MMRole: 1) Instruction Adherence (IA), 2) Fluency (Flu), 3) Coherency (Coh), 4) Image-Text Relevance (ITR), 5) Response Accuracy (RA), 6) Personality Consistency (OC), 7) Knowledge Consistency (KC), and 8) Tone Consistency (TC). On PCogAlignBench, we evaluate: 1) Role-Set Awareness (RSA), 2) Body Behavior Awareness (BBA), 3) Mind Feelings Awareness (MFA), 4) Contextual Awareness (CA), and 5) Conversational Flow (CF). We present the comparison results in Fig. 4, with detailed results provided in Appendix C.2.
@@ 4.4 Analysis
Benefiting from the reduced action space, the constructed latent action space is expected to improve the agent’s rollout diversity during RL exploration, i.e., generating more diverse responses. Prior work has shown that such diversity is critical for improving the upper bound of RL performance li-2025-preserving; yu-2025-dapo.
Following jia-2025-controlling, we quantify rollout diversity via semantic diversity, as it reflects both linguistic diversity and response quality. Concretely, as shown in Fig. 3, for each prompt $(x^{T},x^{T_{1:p}})$ in the RL training set $\mathcal{D}_{\text{RL}}$, the agent generates $G$ responses $\{x^{T_{p+1:m,i}}\}_{i=1}^{G}$, with $p$ as the prompt length and $m$ as the maximum length. We calculate the semantic diversity as:
@@
where $\text{Sim}(\cdot,\cdot)$ denotes the embedding similarity between two responses and we adopt BGE-M3 chen-2024-m3 as the embedding model. We report the average semantic diversity over all samples in the RL training set.
@@
GRPO (Token) 1.079±0.230 1.043±0.197
GRPO (Latent Action) 1.256±0.302 1.200±0.326
Dr.GRPO (Token) 1.074±0.224 1.259±0.249
Dr.GRPO (Latent Action) 1.252±0.297 1.323±0.278
DAPO (Token) 1.075±0.230 1.040±0.182
DAPO (Latent Action) 1.254±0.302 1.131±0.253
BNPO (Token) 1.078±0.224 1.261±0.257
BNPO (Latent Action) 1.297±0.303 1.321±0.286

### v2

@@ Abstract
Vision-language models are increasingly employed as multimodal conversational agents (MCAs) for diverse conversational tasks. Recently, reinforcement learning (RL) has been widely explored for adapting MCAs to various human-AI interaction scenarios. Despite showing great enhancement in generalization performance, fine-tuning MCAs via RL still faces challenges in handling the extremely large text token space. To address this, we learn a compact latent action space for RL fine-tuning instead. Specifically, we adopt the learning from observation mechanism to construct the codebook for the latent action space, where future observations are leveraged to estimate current latent actions that could further be used to reconstruct future observations. However, the scarcity of paired image-text data hinders learning a codebook with sufficient coverage. Thus, we leverage both paired image-text data and text-only data to construct the latent action space, using a cross-modal projector for transforming text embeddings into image-text embeddings. We initialize the cross-modal projector on paired image-text data, and further train it on massive text-only data with a novel cycle consistency loss to enhance its robustness. We show that our latent action based method outperforms competitive baselines on two conversation tasks across various RL algorithms. Code and data are available at https://github.com/AlibabaResearch/DAMO-ConvAI/tree/main/MMLatentAction.
@@ 1 Introduction
Vision-language models (VLMs) Yin et al. (2024) like Qwen-VL Bai et al. (2025a) and GPT-4o Hurst et al. (2024) are increasingly employed as multimodal conversational agents (MCAs) for various conversation tasks Yao et al. (2025). MCAs enable emotionally rich and contextually grounded dialogues based on understanding both input images and texts, and thus become particularly valuable in fields like entertainment Mehta et al. (2022), online education Griol et al. (2014), and personalized assistants Nguyen et al. (2024).
@@
Recently, reinforcement learning (RL) Sutton et al. (1998) has been widely explored for adapting MCAs to diverse real-world human-AI interaction scenarios Zhou et al. (2025). Generally, RL algorithms frame response token generation in MCAs as a sequential decision-making process Chen et al. (2021), which optimize the policy to maximize cumulative rewards through interacting with environments. Despite showing great enhancement in generalization performance Chu et al. (2025), fine-tuning MCAs via RL still faces challenges in dealing with large exploration spaces. For instance, with token vocabulary size $|\mathcal{V}|$ and maximum response length $m$, the sampling space for RL scales exponentially as $|\mathcal{V}|^{m}$.
@@
To address the challenge of large text token space, we learn a compact latent action space for RL fine-tuning instead, following previous works Jia et al. (2025). Specifically, we adopt the learning from observation mechanism Jiang et al. (2023); Ye et al. (2025) to construct the codebook for the latent action space, where future observations are leveraged to estimate current latent actions that could be further used to reconstruct future observations. As a result, the action sampling space at each step is reduced from the token vocabulary size $|\mathcal{V}|$ (e.g., 152K for Qwen2.5-VL Bai et al. (2025b)) to the latent action codebook size $|\mathcal{C}|$ (e.g., 128).
@@
Generally, the codebook has to be learned from diverse data with sufficient coverage, which is a prerequisite for effective RL exploration in latent spaces Chen et al. (2025). Note that VLMs in MCAs are typically pre-trained on paired image-text corpora $(V,T)$, which implicitly convey complementary and partially redundant information between visual and textual modalities Radford et al. (2021). Unfortunately, while unpaired image collections and text corpora are abundant on the web, curating them into aligned image-text corpora remains prohibitively costly Gupta et al. (2025), posing a dilemma in constructing latent spaces. On one hand, using limited paired data and abundant unpaired data would introduce unimodal bias Zhang et al. (2024a), where a model would overly rely on one modality and ignore others. On the other hand, training the codebook solely on limited paired data may result in insufficient coverage, thereby impairing the agent’s generalization ability when handling diverse unseen conversation scenarios.
@@
In this paper, we leverage both paired image-text data $(V,T)$ and unpaired text-only data $T$ to learn the codebook for the latent space. To improve the coverage of latent actions while avoiding potentially unimodal bias, we attempt to construct pseudo paired data $(V^{\prime},T)$ based on text-only data $T$, and use the pseudo data $(V^{\prime},T)$ and the collected data $(V,T)$ to learn the codebook.
@@
However, training a conditional image generator $G(V|T)$ for this purpose is computationally expensive due to the high dimension nature of images Pope et al. (2021). Thus, we learn a cross-modal projector $P$ instead, which transforms an input text $e^{T}$ to an image-text pair $e^{V,T}$ in the embedding space, based on the cross-modal redundancy assumption Radford et al. (2021). Concretely, for each item in the paired image-text data $(V,T)$, we compute the text embedding $e^{T}$ and image-text embedding $e^{V,T}$ using an existing encoder, and train the projector $P$ to imitate the projection between these two kinds of embeddings. To enhance the robustness of the projector $P$, we further train it on massive text-only data $T$ using a cycle consistency loss Zhu et al. (2017). We introduce an additional projector $P^{\prime}$ that can transform image-text embedding $e^{V,T}$ back to text embedding $e^{T}$. In this way, we can optimize the projector $P$ by enforcing cycle consistency on text-only data $T$ such that $P^{\prime}(P(e^{T}))\approx e^{T}$.
@@
We evaluate our method on two conversation tasks, namely multimodal role-playing conversation Dai et al. (2025) and multimodal personalized conversation Li et al. (2025a). To evaluate the generalizability of latent actions, we conduct experiments using various RL algorithms, such as GRPO Shao et al. (2024) and Dr.GRPO Liu et al. (2025). We construct the latent action space using paired image-text data $(V,T)$ and text-only data $T$. The $(V,T)$ data are comprised of image-caption pairs, multimodal news articles, and multimodal Wikipedia pages, totaling 14M images and 1B text tokens. The text-only data are mainly derived from SlimPajama Soboleva et al. (2023), which contains 627B text tokens. Experimental results show that our method outperforms competitive baselines.
@@ 2 Preliminary
In reinforcement learning (RL), problems are framed by a Markov Decision Process (MDP) $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R}\rangle$. For VLMs, the state at step $t$ is the contextual information $s_{t}=(x^{V},x^{T_{1:t}})\in\mathcal{S}$, which includes the input image $x^{V}$ and the current token sequence $x^{T_{1:t}}$. $\mathcal{A}$ is the action space containing all possible actions $a_{t}$ at each step. $\mathcal{T}$ is the state transition function, governing the transition from $s_{t}$ to $s_{t+1}$, i.e., $P\big(s_{t+1}\mid s_{t},a_{t}\big)$. The reward function $\mathcal{R}\big(x^{T_{p+1:m}}\big)$ assigns a scalar reward to the response $x^{T_{p+1:m}}$, conditioned on the input $(x^{V},x^{T_{1:p}})$, with prompt length $p$ and maximum sequence length $m$, following common practice in RL for VLMs Shen et al. (2025).
@@
In traditional token-level RL, each action $a_{t}$ corresponds to selecting the next text token $x^{T_{t+1}}$ from the token vocabulary $\mathcal{V}$, i.e., $\mathcal{A}=\mathcal{V}$. While in latent action RL, at each step $t$, the policy $\pi_{\theta}(a_{t}|x^{V},x^{T_{1:t}})$ selects a latent action $a_{t}$ from a compact codebook $\mathcal{C}$, i.e., $\mathcal{A}=\mathcal{C}$. During RL exploration, the latent action policy samples a latent action at each step, ultimately yielding the terminal state $s_{m}$. During exploitation, the latent action policy is refined to maximize expected rewards using RL algorithms such as GRPO Shao et al. (2024).
@@ 3.2 Latent Action Space Learning
Following Jia et al. (2025), we construct the latent action space using large-scale corpora in two steps. 1) inverse dynamics learning, which trains the $f_{\text{world}}$, $f_{\text{inverse}}$, and $\mathcal{C}$ in an unsupervised manner (Fig. 2 (a)); 2) policy behavior cloning, which trains the policy model $\pi_{\theta}$ to mimic the latent action $a_{t}$ inferred by $f_{\text{inverse}}$ (Fig. 2 (b)).
@@ 3.3 Latent Action Reinforcement Learning
where $R(\cdot)$ denotes the reward function. During latent action RL, we only optimize the latent action prediction distribution of the policy model while keeping the language world model (responsible for text token generation) frozen. Consequently, the length of the optimized sequence (latent action sequence) is equal to the response length, which means that our approach does not increase the length of the optimized sequence compared to token-level RL methods.
@@ 4.1 Experimental Setup
We build the language world model, inverse dynamics model, and policy model upon the same foundation vision-language model. Specifically, we use Qwen2.5-VL-3B-Instruct and Qwen2.5-VL-7B-Instruct Bai et al. (2025b) for main experiments. The latent action space is implemented as a codebook with size $|\mathcal{C}|=128$.
@@
During the latent action space construction stage (Section 3.2), we use a mixture of paired image-text corpora $\mathcal{D}^{VT}$ and text-only corpora $\mathcal{D}^{T}$. For $\mathcal{D}^{VT}$, we collect image-caption pairs from Conceptual-12M Changpinyo et al. (2021), multimodal news articles from N24News Wang et al. (2022), and multimodal Wikipedia data from WikiWeb2M Burns et al. (2023), totaling 14 million images and 1 billion text tokens. For $\mathcal{D}^{T}$, we collect text-only data mainly from the SlimPajama-627B dataset Soboleva et al. (2023), which contains 627 billion text tokens.
For latent action RL (Section 3.3), we evaluate our method on two downstream tasks: 1) multimodal role-playing conversation on MMRole Dai et al. (2025), where we focus on the challenging Comment subset; we train on the in-distribution (ID) split and evaluate on ID and out-of-distribution (OOD) test sets; 2) multimodal personalized conversation on PCogAlignBench Li et al. (2025a), where we train the agent on the LS1 set and evaluate on LS1 and LS2 test sets.
We select MMRole and PCogAlignBench because their responses average $\approx 200$ tokens (Figure 6 and 7), yielding a significantly larger RL sampling space than traditional conversational benchmarks Haber et al. (2019); De Vries et al. (2017); Das et al. (2017), whose responses are typically less than 20 tokens. Since the RL sampling space grows exponentially with response length Feng et al. (2025), these benchmarks pose substantially greater challenges for RL fine-tuning, making them more suitable for evaluating our method.
@@
We adopt the LLM-as-a-Judge metric to evaluate model performance, using prompt templates validated by Dai et al. (2025); Li et al. (2025a), which show high correlation with human judgments. For each sample, the LLM judge scores both the model and ground-truth responses across benchmark-specific dimensions, with scores ranging 1-10. Then, following Dai et al. (2025), we report the ratio of the model’s average score to the ground-truth response’s average score across all evaluation dimensions. We report the mean and standard deviation across three evaluation runs.
@@
We consider two categories of baselines: 1) Non-RL baselines: the naive Prompt and supervised fine-tuning (SFT); 2) RL-based methods, where we compare two optimization strategies, token-level and latent action RL, using four algorithms: a) Group Relative Policy Optimization (GRPOShao et al. (2024), b) Dr. GRPO Liu et al. (2025), c) Decoupled Clip and Dynamic Sampling Policy Optimization (DAPOYu et al. (2025), and d) Beta Normalization Policy Optimization (BNPOXiao et al. (2025). The reward functions are kept the same for methods. Please refer to the Appendix B for more experimental details.
@@ 4.2 Main Results
To thoroughly evaluate the performance of multimodal conversational agents trained with latent actions across various fine-grained conversational dimensions, following prior work Dai et al. (2025); Li et al. (2025a), we assess eight dimensions on MMRole: 1) Instruction Adherence (IA), 2) Fluency (Flu), 3) Coherency (Coh), 4) Image-Text Relevance (ITR), 5) Response Accuracy (RA), 6) Personality Consistency (OC), 7) Knowledge Consistency (KC), and 8) Tone Consistency (TC). On PCogAlignBench, we evaluate: 1) Role-Set Awareness (RSA), 2) Body Behavior Awareness (BBA), 3) Mind Feelings Awareness (MFA), 4) Contextual Awareness (CA), and 5) Conversational Flow (CF). We present the comparison results in Fig. 4, with detailed results provided in Appendix C.3.
@@ 4.4 Analysis
Benefiting from the reduced action space, the constructed latent action space is expected to improve the agent’s rollout diversity during RL exploration, i.e., generating more diverse responses. Prior work has shown that such diversity is critical for improving the upper bound of RL performance Li et al. (2025b); Yu et al. (2025).
Following Jia et al. (2025), we quantify rollout diversity via semantic diversity, as it reflects both linguistic diversity and response quality. Concretely, as shown in Fig. 3, for each prompt $(x^{T},x^{T_{1:p}})$ in the RL training set $\mathcal{D}_{\text{RL}}$, the agent generates $G$ responses $\{x^{T_{p+1:m,i}}\}_{i=1}^{G}$, with $p$ as the prompt length and $m$ as the maximum length. We calculate the semantic diversity as:
@@
where $\text{Sim}(\cdot,\cdot)$ denotes the embedding similarity between two responses and we adopt BGE-M3 Chen et al. (2024) as the embedding model. We report the mean and standard deviation of the semantic diversity computed over 5 independent runs, where the standard deviation reflects the variability of rollout diversity across different seeds.
@@
GRPO (Token) 1.079±0.001 1.042±0.001
GRPO (Latent Action) 1.248±0.002 1.191±0.002
Dr.GRPO (Token) 1.070±0.001 1.256±0.002
Dr.GRPO (Latent Action) 1.246±0.001 1.318±0.002
DAPO (Token) 1.073±0.001 1.038±0.001
DAPO (Latent Action) 1.253±0.001 1.127±0.001
BNPO (Token) 1.077±0.001 1.257±0.003
BNPO (Latent Action) 1.291±0.002 1.315±0.002